# Classification of abelian Schur groups I

### v1

@@ 7.3. Dense $S$-rings
One can see that $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$ for some $K_{H}\leq\operatorname{Aut}(H)$ by Lemma 7.8. If $\mathcal{A}_{U}=\operatorname{Cyc}(K_{U},U)$ for some $K_{U}\leq\operatorname{Aut}(U)$, then $\mathcal{A}_{V}=\operatorname{Cyc}(K_{H}\times K_{U},V)$ by Eqs. (4) and (31). Therefore Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Otherwise, $\mathcal{A}_{U}=\mathcal{T}_{U}$ by Lemma 6.2. This yields that $L=U$. Together with Eq. (31), this implies that $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{V}$. Thus, Statement $(2)$ from Theorem 7.1 holds for $S$ and we are done.
@@ 7.5. Auxiliary statements
Let $n=p^{k}$, $p\geq 5$, and $S$ an $\mathcal{A}$-section such that $p^{2}~|~|S|$, $\mathcal{A}_{S}$ is indecomposable, and there is an $\mathcal{A}_{S}$-subgroup of order or index $p$. Then $\mathcal{A}_{S}$ is normal.
@@
Clearly, $\mathcal{A}_{S}$ is nontrivial. Firstly, suppose that $S$ is cyclic. If $|S|$ is a power of $p$, then $\mathcal{A}_{S}$ is normal by Lemma 6.2. If $|S|=2p^{l}$ for some $l\geq 2$ and $\mathcal{A}_{S}$ is not normal, then $\mathcal{A}_{S}=\mathbb{Z}S_{1}\otimes\mathcal{T}_{S_{2}}$, where $S_{1},S_{2}\leq S$ with $|S_{1}|=2$ and $|S_{2}|=p^{l}$, by Lemma 6.13. However, the latter equality contradicts to the condition of the lemma that there is an $\mathcal{A}_{S}$-subgroup of order or index $p$.
Now suppose that $S$ is noncyclic, i.e. 4||S|4~|~|S|. Let
$\mathcal{A}_{S}=\mathcal{A}_{S_{1}}\otimes\mathcal{A}_{S_{2}}$
for some proper nontrivial $\mathcal{A}_{S}$-subgroups $S_{1}$ and $S_{2}$ such that $S=S_{1}\times S_{2}$. Without loss of generality, we may assume that $|S_{1}|=4$ and $|S_{2}|=p^{l}$ or $|S_{1}|=2$ and $|S_{2}|=2p^{l}$ for some $l\geq 2$. In particular, $S_{2}$ is cyclic. Note that $|\operatorname{rad}(\mathcal{A}_{S_{2}})|=1$ because otherwise $\mathcal{A}_{S_{2}}$ is decomposable by Lemma 6.2 and hence $\mathcal{A}_{S}$ so is. If $|S_{1}|=4$ ($|S_{1}|=2$, respectively), then the last condition of Lemma 7.14 and Lemma 6.2 (Lemma 6.13, respectively) imply that $\mathcal{A}_{S_{2}}$ is normal. One can see that $\mathcal{A}_{S_{1}}$ is normal. Indeed, this is clear if $|S_{2}|=2$ and follows from Lemma 7.8 if $|S_{2}|=4$. Therefore $\mathcal{A}_{S}$ is normal by Eq. (3).
In view of the above paragraph, we may assume that $\mathcal{A}_{S}$ is not a nontrivial tensor product. Since $\mathcal{A}_{S}$ is also nontrivial and indecomposable, Propositions 7.3 and 7.4 can not hold for $\mathcal{A}$. So $\mathcal{A}_{S}$ is dense. In particular, the Sylow $p$-subgroup $S_{2}$ of $S$ is an $\mathcal{A}_{S}$-subgroup. If $|\operatorname{rad}(\mathcal{A}_{S_{2}})|>1$, then $\mathcal{A}_{S}$ is decomposable or $p=3$ by Proposition 7.12, a contradiction to the assumption of the lemma. Therefore $|\operatorname{rad}(\mathcal{A}_{S_{2}})|=1$. Thus, $\mathcal{A}_{S}$ is normal by Proposition 7.11(1). ∎
@@
Let $n=p^{k}$ and $\mathcal{A}$ an $S$-ring over $G$. The following statements hold.
@@
If $p=3$, then $\operatorname{Hol}(C_{3})\cong S_{3}$ and the lemma is clear. Let $p\geq 5$. Let us prove Statement $(1)$. If $P\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$, then $\mathcal{A}$ is the $U/P$-wreath product for an $\mathcal{A}$-subgroup $U\leq H\times P$ and we are done. Further, we assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$ such that
@@
Let $R=\operatorname{rad}(X)$ and $V=\langle X\rangle$. Clearly, $R\leq H$. As $|R\cap P|=1$, the kernel of the action of $\operatorname{Aut}(\mathcal{A})^{P}$ on $P/R$ is trivial and hence it is enough to show that
$\operatorname{Aut}(\mathcal{A})^{P/R}\leq\operatorname{Hol}(P/R).$
Due to the definitions of $R$ and $V$, we conclude that $\mathcal{A}_{V/R}$ is indecomposable. Clearly, $P/R$ is an $\mathcal{A}_{V/R}$-subgroup of order $p$. One can see that $p^{2}~|~|V/R|$ because $X\nsubseteq H\times P$ and $R\leq H$. So $V/R$ satisfies all the conditions of Lemma 7.14 and consequently $\mathcal{A}_{V/R}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{P/R}\leq\operatorname{Aut}(\mathcal{A}^{V/R})^{P/R}\leq\operatorname{Aut}(\mathcal{A}_{V/R})^{P/R}\leq(\operatorname{Hol}(V/R))^{P/R}\leq\operatorname{Hol}(P/R),$
where the third inclusion follows from the normality of $\mathcal{A}_{V/R}$, and we are done.
@@
Let $V=\langle X\rangle$. Clearly, $|G:V|\in\{1,2,4\}$ and $|V:(V\cap U)|=|G:U|=p$. Observe that $\operatorname{Aut}(\mathcal{A})^{G/U}$ and $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ are isomorphic as permutation groups. Indeed, the kernel of the action of $\operatorname{Aut}(\mathcal{A})$ on $G/U$ contains $H_{r}$ and hence for every $f\in\operatorname{Aut}(\mathcal{A})$, there is $f^{\prime}\in\operatorname{Aut}(\mathcal{A})$ such that $f^{G/U}=(f^{\prime})^{G/U}$ and $V^{f^{\prime}}=V$. It is easy to see that for every $\Delta\in G/U$, there is a unique $\Delta^{\prime}\in V/(V\cap U)$ such that $\Delta^{\prime}\subseteq\Delta$. Therefore the mappings $f^{G/U}\mapsto(f^{\prime})^{V/(V\cap U)}$ from $\operatorname{Aut}(\mathcal{A})^{G/U}$ to $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ and $\Delta\mapsto\Delta^{\prime}$ provide an isomorphism of permutation groups $\operatorname{Aut}(\mathcal{A})^{G/U}$ and $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$. Thus, it is enough to show that
@@
Let $R=\operatorname{rad}(X)$. Clearly, the $S$-ring $\mathcal{A}_{V/R}$ is indecomposable and $(V\cap U)/R$ is an $\mathcal{A}_{V/R}$-subgroup of index $p$. Since $R\ngeq(U\cap D)$, we conclude that $p^{2}~|~|V/R|$. So $V/R$ satisfies all the conditions of Lemma 7.14 and hence $\mathcal{A}_{V/R}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}^{V/R})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}_{V/R})^{V/(V\cap U)}\leq$
$\leq(\operatorname{Hol}(V/R))^{V/(V\cap U)}\leq\operatorname{Hol}(V/(V\cap U)),$
where the third inclusion follows from the normality of $\mathcal{A}_{V/R}$, and we are done. ∎
It should be mentioned that if $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) in Statement $(1)$ (Statement $(2)$, respectively) of Lemma 7.15 is nontrivial, then $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) is normal by Lemma 6.2 and the claim of Lemma 7.15 easily follows from the fact that an automorphism group of a normal $S$-ring is a subgroup of a holomorph of the underlying group. We also note that if the generalized wreath product in Statement $(1)$ or $(2)$ of Lemma 7.15 is trivial, then $|G|=4p$.
@@
Throughout this section, we keep the notation from the previous one. We start with a lemma concerned with a schurity of proper subgroups of $G$.
@@
Let $N<G$. If $N$ is cyclic, then $N$ is Schur by [9, Theorem 1.1], whereas if $N\cong E_{4}$, then $N$ is Schur by [10, Theorem 1.2]. Otherwise, $N\cong E_{4}\times C_{r^{m}}$, where $r\in\{p,q\}$ and $1\geq m<k$, is Schur by induction on $m$ whose base follows from [10, Theorem 1.5]. ∎
@@
One can see that $\mathcal{A}_{S_{D}}$ is cyclotomic because $\mathcal{A}_{W}$ so is. Suppose that Statement $(3)$ of Theorem 7.1 holds. Then $|\operatorname{rad}(\mathcal{A}_{W})|=1$ and $|W_{H}|\leq 2$. Observe that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$. Indeed, Lemma 3.2(2) and Lemma 6.2 imply that $\mathcal{A}=\mathbb{Z}W_{H}\otimes\mathcal{A}_{W_{D}}$ and hence $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=|\operatorname{rad}(\mathcal{A}_{W})|=1$. If $S_{D}=W_{D}$, then obviously $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$. Otherwise, $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 6.9.
@@
If $|S_{H}|\leq 2$, then $\mathcal{A}_{S_{H}}=\mathbb{Z}S_{H}$ and Eq. (32) follows from Lemma 3.2(2). If $|S_{H}|=4$ and Eq. (32) does not hold, then $\mathcal{A}_{S}$ is $2$-minimal by Proposition 7.11(1) and Lemma 8.2, a contradiction to the assumption.
@@
Now suppose that $|V|=4p$. If $\mathcal{A}_{V}$ is dense, then $\mathcal{A}$ is cyclotomic by [10, Case 33, p. 115]. Therefore $\mathcal{A}_{V}=V(K_{0},V)$ for some $K_{0}\leq\operatorname{Hol}(V)$. Clearly, $K_{0}^{S}\leq\operatorname{Hol}(S)$ and $\mathcal{A}_{S}=V(K_{0}^{S},S)$. Since $\mathcal{A}_{S}$ is Cayley minimal (Lemma 6.4) and $\mathcal{A}_{S}=V(\operatorname{Hol}(S),S)$, we conclude that $K_{0}^{S}=\operatorname{Hol}(S)$. Together with Eq. (33) and Lemma 3.5, this implies that $\mathcal{A}$ is schurian.
Let $\mathcal{A}$ be nondense. If $V=G/L$, then $\mathcal{A}_{V}=\mathcal{A}_{G/L}$ is the $S/S_{D}$-wreath product by [10, Case 11, p. 114] and hence $\mathcal{A}$ is the $S_{1}=U/L_{1}$-wreath product, where $L_{1}=(S_{D})^{\pi^{-1}}$. If $V=U$, then $\mathcal{A}_{V}=\mathcal{A}_{U}$ is the $S_{2}=H/L$-wreath product by [10, Case 22, p. 114] (in this case, $|L|=2$) and hence $\mathcal{A}$ is the $S_{2}$-wreath product. One can see that $|S_{1}|=|U/L_{1}|=|S/S_{D}|=|S_{H}|=2$ and $|S_{2}|=|H/L|=2$. Thus, $\mathcal{A}$ is schurian by Lemma 3.6. ∎
@@
Further, we assume that $|S_{H}|=4$ and $\mathcal{A}_{S_{H}}=\mathcal{T}_{S_{H}}$. Due to Remark 2.7, there is a unique proper $2$-equivalent subgroup $K_{\min}$ of $\operatorname{Aut}(\mathcal{A}_{S_{H}})$ of order $12$ such that $K_{\min}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}_{U}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U\cap D}$, then put $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$. In this case, $K_{1}^{S_{H}}=K_{\min}$ by Lemma 7.13. Otherwise, $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{H})\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$. Let $K_{H}\leq\operatorname{Sym}(H)$ such that $K_{H}\cong K_{\min}$. The group $K_{H}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{H})$ (see Remark 2.7) and hence $K_{1}=K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{U})$. By the definition of $K_{1}$,
$K_{1}^{S_{H}}=(K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D}))^{S_{H}}=(K_{H})^{S_{H}}=K_{\min}.$
The groups $S_{H}$ and $D/L$ are $\mathcal{A}_{G/L}$-subgroups. If $\mathcal{A}_{G/L}\neq\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{D/L}$, then put $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$; otherwise put $K_{0}=K_{\min}\times\operatorname{Aut}(\mathcal{A}_{D/L})$. The argument which is the same to the argument from the previous paragraph implies that $K_{0}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{G/L})$ and $K_{0}^{S_{H}}=K_{\min}=K_{1}^{S_{H}}$ as required. ∎

### v3

@@ 7.3. Dense $S$-rings
Then $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$ for some $K_{H}\leq\operatorname{Aut}(H)$ by Lemma 7.9. If $\mathcal{A}_{U}=\operatorname{Cyc}(K_{U},U)$ for some $K_{U}\leq\operatorname{Aut}(U)$, then $\mathcal{A}_{V}=\operatorname{Cyc}(K_{H}\times K_{U},V)$ by Eqs. (4) and (35). Therefore Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Otherwise, $\mathcal{A}_{U}=\mathcal{T}_{U}$ by Lemma 6.2. Then $L=U$. Together with Eq. (35), this implies that $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{V}$. Thus, Statement $(2)$ from Theorem 7.1 holds for $S$ and we are done.
@@ 7.5. Auxiliary statements
Let $\mathcal{B}$ be an $S$-ring over a subgroup $V$ of $E_{4}\times C_{p^{k}}$, where $p\geq 5$ and $k\geq 2$. Suppose that $p^{2}~|~|V|$, $\mathcal{B}$ is indecomposable, and there is a $\mathcal{B}$-subgroup of order or index $p$. Then $\mathcal{B}$ is normal.
@@
Clearly, $\mathcal{B}$ is nontrivial. Firstly, suppose that $V$ is cyclic. If $|V|$ is a $p$-power, then $\mathcal{B}$ is normal by the supposition of this lemma and Lemma 6.2. If $|V|=2p^{l}$ for some $l\geq 2$ and $\mathcal{B}$ is not normal, then $\mathcal{B}=\mathcal{T}_{V_{1}}\otimes\mathbb{Z}V_{2}$, where $V_{1},V_{2}\leq V$ with $|V_{1}|=p^{l}$ and $|V_{2}|=2$, by Lemma 6.13. However, the latter equality contradicts to the supposition of the lemma that there is a $\mathcal{B}$-subgroup of order or index $p$.
Now suppose that $V$ is noncyclic, i.e. 4||V|4~|~|V|. Let
$\mathcal{B}=\mathcal{B}_{V_{1}}\otimes\mathcal{B}_{V_{2}}$
for some proper nontrivial $\mathcal{B}$-subgroups $V_{1}$ and $V_{2}$ such that $V=V_{1}\times V_{2}$. Without loss of generality, we may assume that $|V_{1}|=p^{l}$ and $|V_{2}|=4$, or $|V_{1}|=2p^{l}$ and $|V_{2}|=2$ for some $l\geq 2$. In particular, $V_{1}$ is cyclic. Note that $|\operatorname{rad}(\mathcal{B}_{V_{1}})|=1$ because otherwise $\mathcal{B}_{V_{1}}$ is decomposable by Lemma 6.2 and hence so is $\mathcal{B}$, which contradicts to the supposition of the lemma. If $|V_{1}|=p^{l}$ ($|V_{2}|=2p^{l}$, respectively), then the last condition of this lemma and Lemma 6.2 (Lemma 6.13, respectively) imply that $\mathcal{B}_{V_{1}}$ is normal. Then $\mathcal{B}_{V_{2}}$ is also normal. Indeed, this is clear if $|V_{2}|=2$ and follows from Lemma 7.9 if $|V_{2}|=4$. Therefore $\mathcal{B}$ is normal by Eq. (3).
In view of the above paragraph, we may assume that $\mathcal{B}$ is not a nontrivial tensor product. Since $\mathcal{B}$ is also nontrivial and indecomposable, Theorems 7.3 and 7.8 can not hold for $\mathcal{B}$. So $\mathcal{B}$ is dense. In particular, the Sylow $p$-subgroup $V_{1}$ of $S$ is a $\mathcal{B}$-subgroup. If $|\operatorname{rad}(\mathcal{B}_{V_{1}})|>1$, then $\mathcal{B}$ is decomposable or $p=3$ by Theorem 7.13, a contradiction to the supposition of the lemma. Therefore $|\operatorname{rad}(\mathcal{B}_{V_{1}})|=1$. Thus, $\mathcal{B}$ is normal by Theorem 7.12(1). ∎
Recall that $\mathcal{A}$ is an $S$-ring over $G=H\times D$, where $H\cong E_{4}$ and $D\cong C_{n}$ for $n\in\{pq,p^{k}\}$.
@@
Let $n=p^{k}$. The following statements hold.
@@
If $p=3$, then $\operatorname{Hol}(C_{3})\cong\operatorname{Sym}(3)$ and the lemma is clear. Let $p\geq 5$. Let us prove Statement $(1)$. If $P\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$, then $\mathcal{A}$ is the $U/P$-wreath product for an $\mathcal{A}$-subgroup $U\leq H\times P$ and we are done. Further, we assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$ such that
@@
Let $R=\operatorname{rad}(X)$ and $V=\langle X\rangle$. The image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/R$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. As $R\ngeq P$, we have $R\leq H$ and $|R\cap P|=1$ (note that $R$ can be trivial). The latter equality implies that the kernel of the action of $\operatorname{Aut}(\mathcal{A})^{P}$ on $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ is trivial and hence it suffices to show that
$\operatorname{Aut}(\mathcal{A})^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu).$
Since $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$ generates $\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu$ and $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1$, the $S$-ring $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable. Clearly, $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ is an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$-subgroup of order $p$. Moreover, $p^{2}~|~|\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu|$ because $X\nsubseteq H\times P$ and $R\leq H$. So $\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu$ satisfies all the conditions of Lemma 7.15 and consequently $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq(\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu))^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu),$
where the second inequality follows from the normality of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$, and we are done.
@@
Let $V=\langle X\rangle$. Clearly, $|G:V|\in\{1,2,4\}$ and $|V:(V\cap U)|=|G:U|=p$. Observe that $\operatorname{Aut}(\mathcal{A})^{G/U}$ and $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ are permutation isomorphic in the sense of [4, p. 17]. Indeed, the kernel of the action of $\operatorname{Aut}(\mathcal{A})$ on $G/U$ contains $H_{r}$ and hence for every $f\in\operatorname{Aut}(\mathcal{A})$, there is $f^{\prime}\in\operatorname{Aut}(\mathcal{A})$ such that $f^{G/U}=(f^{\prime})^{G/U}$ and $V^{f^{\prime}}=V$. As $V\cap U\leq U$, for every $\Delta\in G/U$, there is $\Delta^{\prime}\in V/(V\cap U)$ such that $\Delta^{\prime}\subseteq\Delta$. Moreover, such $\Delta^{\prime}$ is unique because $|V:(V\cap U)|=|G:U|=p$. Therefore the mappings $f^{G/U}\mapsto(f^{\prime})^{V/(V\cap U)}$ from $\operatorname{Aut}(\mathcal{A})^{G/U}$ to $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ and $\Delta\mapsto\Delta^{\prime}$ from $G/U$ to $V/(V\cap U)$ define a permutation isomorphism from $\operatorname{Aut}(\mathcal{A})^{G/U}$ to $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$. Thus, it is enough to show that
@@
Let $R=\operatorname{rad}(X)$. Again, the image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/R$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. Clearly, the $S$-ring $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable and $\mkern 1.5mu\overline{\mkern-1.5muV\cap U\mkern-1.5mu}\mkern 1.5mu$ is an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$-subgroup of index $p$. Since $R\ngeq(U\cap D)$, we conclude that $p^{2}~|~|\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu|$. So $\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu$ satisfies all the conditions of Lemma 7.15 and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}^{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu})^{V/(V\cap U)}\leq$
$\leq(\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu))^{V/(V\cap U)}\leq\operatorname{Hol}(V/(V\cap U)),$
where the third inequality follows from the normality of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$, and we are done. ∎
It should be mentioned that if $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) in Statement $(1)$ (Statement $(2)$, respectively) of Lemma 7.16 is nontrivial, then $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) is normal by Lemma 6.2 because $P$ ($G/U$, respectively) is of prime order $p$. Therefore the claim of Lemma 7.16 easily follows from the definition of a normal $S$-ring. We also note that if the generalized wreath product in Statement $(1)$ or $(2)$ of Lemma 7.16 is trivial, then $|G|=4p$.
@@
Throughout this section, we keep the notation from Section 7. We start with a lemma concerned with a schurity of proper subgroups of $G$.
@@
Let $N<G$. If $N$ is cyclic, then $N$ is Schur by [10, Theorem 1.1], whereas if $N\cong E_{4}$, then $N$ is Schur by [11, Theorem 1.2]. Otherwise, $N\cong E_{4}\times C_{r^{m}}$, where $r\in\{p,q\}$ and $k>m\geq 1$, is Schur by induction on $m$ whose base follows from [11, Theorem 1.5]. ∎
@@
One can see that $\mathcal{A}_{S_{D}}$ is cyclotomic because so is $\mathcal{A}_{W}$. Suppose that Statement $(3)$ of Theorem 7.1 holds. Then $|\operatorname{rad}(\mathcal{A}_{W})|=1$ and $|W_{H}|\leq 2$. Observe that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$. Indeed, Lemma 3.2(2) and Lemma 6.2 imply that $\mathcal{A}_{W}=\mathbb{Z}W_{H}\otimes\mathcal{A}_{W_{D}}$ and hence $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=|\operatorname{rad}(\mathcal{A}_{W})|=1$. If $S_{D}=W_{D}$, then obviously $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$. Otherwise, $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 6.9.
@@
If $|S_{H}|\leq 2$, then $\mathcal{A}_{S_{H}}=\mathbb{Z}S_{H}$ and Eq. (36) follows from Lemma 3.2(2). If $|S_{H}|=4$ and Eq. (36) does not hold, then $\mathcal{A}_{S}$ is $2$-minimal by Theorem 7.12(1) and Lemma 8.2, a contradiction to the assumption.
@@
Now suppose that $|V|=4p$. If $\mathcal{A}_{V}$ is dense, then $\mathcal{A}$ is cyclotomic by [11, Case 33, p. 115]. Therefore $\mathcal{A}_{V}=V(K_{0},V)$ for some $K_{0}\leq\operatorname{Hol}(V)$. Clearly, $K_{0}^{S}\leq\operatorname{Hol}(S)$ and $\mathcal{A}_{S}=V(K_{0}^{S},S)$. Since $\mathcal{A}_{S}$ is Cayley minimal (Lemma 6.4) and $\mathcal{A}_{S}=V(\operatorname{Hol}(S),S)$, we conclude that $K_{0}^{S}=\operatorname{Hol}(S)$. Together with Eq. (37) and Lemma 3.5, this implies that $\mathcal{A}$ is schurian.
Let $\mathcal{A}$ be nondense. If $V=G/L$, then $\mathcal{A}_{V}=\mathcal{A}_{G/L}$ is the $S/S_{D}$-wreath product by [11, Case 11, p. 114] and hence $\mathcal{A}$ is the $S_{1}=U/L_{1}$-wreath product, where $L_{1}=(S_{D})^{\pi^{-1}}$. If $V=U$, then $\mathcal{A}_{V}=\mathcal{A}_{U}$ is the $S_{2}=H/L$-wreath product by [11, Case 22, p. 114] (in this case, $|L|=2$) and hence $\mathcal{A}$ is the $S_{2}$-wreath product. One can see that $|S_{1}|=|U/L_{1}|=|S/S_{D}|=|S_{H}|=2$ and $|S_{2}|=|H/L|=2$. Thus, $\mathcal{A}$ is schurian by Lemma 3.6. ∎
@@
Further, we assume that $|S_{H}|=4$ and $\mathcal{A}_{S_{H}}=\mathcal{T}_{S_{H}}$. By Remark 2.7, there is a unique proper $2$-equivalent subgroup $\operatorname{Alt}(S_{H})\cong\operatorname{Alt}(4)$ of $\operatorname{Aut}(\mathcal{A}_{S_{H}})\cong\operatorname{Sym}(4)$ of order $12$ such that $\operatorname{Alt}(S_{H})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}_{U}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U\cap D}$, then put $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$. In this case, $K_{1}^{S_{H}}=\operatorname{Alt}(S_{H})$ by Lemma 7.14. Otherwise, $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{H})\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$. Let $K_{H}\leq\operatorname{Sym}(H)$ be such that $K_{H}\cong\operatorname{Alt}(S_{H})$. The group $K_{H}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{H})$ (see Remark 2.7) and hence $K_{1}=K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{U})$. By the definition of $K_{1}$,
$K_{1}^{S_{H}}=(K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D}))^{S_{H}}=(K_{H})^{S_{H}}=\operatorname{Alt}(S_{H}).$
The groups $S_{H}$ and $D/L$ are $\mathcal{A}_{G/L}$-subgroups. If $\mathcal{A}_{G/L}\neq\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{D/L}$, then put $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$; otherwise put $K_{0}=\operatorname{Alt}(S_{H})\times\operatorname{Aut}(\mathcal{A}_{D/L})$. The argument which is the similar to the argument from the previous paragraph implies that $K_{0}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{G/L})$ and $K_{0}^{S_{H}}=\operatorname{Alt}(S_{H})=K_{1}^{S_{H}}$ as required. ∎